7. A particle is projected horizontally at 15ms-1 from height of 20m. Calculate the horizontal distance covered
by the particle just before hitting the ground (g = 10ms-2).
1
Expert's answer
2020-04-27T09:56:31-0400

Given:

"{v_0=15\:\rm m/s},\\
{H=20\:\rm m/s},\\
{g=10\:\rm m/s^2}."

The horizontal distance


"H=v_0\sqrt{\frac{2H}{g}}=15\sqrt{\frac{2\times 20}{10}}=30\:\rm m."